OpenStudy (anonymous):

If 36 g of Cobalt (II) chloride is used in the reaction with 24 g of Mercury (II) phosphate, how much Cobalt (II) phosphate can be produced? What is the limiting reactant?

OpenStudy (dtan5457):

This one was a bit tricky, but I gave it a shot anyways. Cobalt chloride=CoCl2 Mercury phosphate=Hg3(PO4)2 Cobalt phosphate=Co3(PO4)2 Your equation is basically CoCl2+Hg3(PO4)2>>>>Co3(PO4)2 From there you make two ratios one for the problem which will be the moles of Cocl2 and the moles of Hg3(Po4)2 that will be 0.28/0.03=9.3 the equation ratio is just 1:1 in this case the problem ratio>equation ratio...so the demoninator is limiting and the numerator is excess excess=Cocl2 limiting=Hg3(PO4)2 now to get the mass for Cobalt (II) phosphate. take the limiting reactant from the problem, multiply it by the excess in the equation, and the molar mass of Cobalt (II) phosphate 0.28 x 1 x366=roughly 102.48 grams of Cobalt (II) phosphate

OpenStudy (anonymous):

That was impressive, thank you, I get it now, I have another one for you if you are up to it. Just watch the feed.

OpenStudy (dtan5457):

I'll try to get to it.